OpenStudy (anonymous):

Fool's problem of the day, (1) If \( p_1, p_2, p_3\) are respectively the perpendicular from the vertices of \( \triangle ABC \) to the opposite sides, then prove that \[ \hspace{3cm} \frac {\cos A}{p_1} + \frac {\cos B}{p_2} + \frac {\cos C}{p_3} = \frac 1 R \] where \( R \) is the circumradius of the \( \triangle ABC \) . Good luck!

OpenStudy (anonymous):

I am unable to draw me a nice diagram :(

OpenStudy (anonymous):

Diagram is not necessary to solve this problem.

OpenStudy (blockcolder):

What's a circumradius?

OpenStudy (anonymous):

@blockcolder: http://mathworld.wolfram.com/Circumradius.html

OpenStudy (henryblah):

I think I saw something really similar to this on an AEA paper the other day...

OpenStudy (anonymous):

Why can't I solve it? What's making this question so hard to solve?

OpenStudy (anonymous):

So far,I got 1/R = (CosA+CosB +CosC-1)/(p_1^-1+p_2^-1+p_3^-1)

OpenStudy (anonymous):

This requires some knowledge of certain other identities.

OpenStudy (anonymous):

Isn't Law of Sines enough? :D

OpenStudy (anonymous):

No

OpenStudy (anonymous):

You need things with R (and maybe inradius as well)

OpenStudy (anonymous):

I used \[\sin 2A + \sin 2B + \sin 2C = 4 \sin A \sin B \sin C \]

OpenStudy (anonymous):

and \[ \large R= \frac {abc}{4 \times Area} \]

OpenStudy (anonymous):

Yes, I was wondering about that last one.

OpenStudy (anonymous):

I was trying to use like bc = 2R p1

OpenStudy (anonymous):

Using the Law of Sines, and Trig. substitution.\[\frac3R = \frac{2\sin c \sin b }{p_1} + \frac{2\sin a \sin c}{p_2}+ \frac{2\sin b \sin a}{p_3}\tag 1\] \[\frac3R = \underbrace{\frac{\cos a}{p_1}+\frac{\cos b}{p_2}+\frac{\cos c}{p_3}}_{x} + \frac{\cos (b-c)}{p_1}+\frac{\cos (a-c)}{p_2}+\frac{\cos (b-a)}{p_3}\tag 2\] \[\small\frac3R = x + \underbrace{\frac{\sin c \sin b }{p_1} + \frac{\sin a \sin c}{p_2}+ \frac{\sin b \sin a}{p_3}}_{\frac{3}{R2},\quad \mathsf{\text{From (1)}}} + \frac{\cos c\cos b}{p_1}+\frac{\cos c\cos a}{p_2}+\frac{\cos b\cos a}{p_3}\] \[\frac{3}{R} = x + \frac{\cos c\cos b}{p_1}+\frac{\cos c\cos a}{p_2}+\frac{\cos b\cos a}{p_3}\tag 3\] Something is wrong with my method, if you will compare (1),(2) and (3). You will see it. \[2R = \frac{a}{\sin a}= \frac{b}{\sin b} = \frac{c}{\sin c} \mathsf{\tag {Law of Sines}}\]  \[a=\frac{p_3}{\sin b}, \quad b =\frac{p_1}{\sin c},\quad c = \frac{p_2}{\sin a}. \]